OpenStudy (charlotte123):

In a certain country ½ of 5 = 3. If the same proportion holds, what is the value of 1/3 of 10 ?

OpenStudy (shamil98):

5/2 = 3 and 10/3 = x (10/3)/(5/2) = 20/15 = 4/3 So x/3 = 4/3 x=4

OpenStudy (charlotte123):

dang it sham =,=

OpenStudy (shamil98):

LEEEEEEEEEEEEEEEEDLE LEEEEEEEEEEEEEEEDLE LEEEEEEEEEEDLE

OpenStudy (charlotte123):

PUMPKIN o-o